OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y=5x+8; (2,16)

OpenStudy (texaschic101):

y = 5x + 8 -- when your equation is in y = mx + b form, the m position is your slope. The slope of this equation is 5. A parallel line will have the same slope. we will now use y = mx + b form slope(m) = 5 (2,16) x = 2 and y = 16 now we sub 16 = 5(2) + b 16 = 10 + b 16 - 10 = b 6 = b so your parallel equation is : y = 5x + 6 any questions ?

OpenStudy (anonymous):

Nope, no questions thank you! :)

OpenStudy (texaschic101):

:)